OpenStudy (anonymous):

Determine the values m, M, a, b, and c such that the function f(x) satisfies the hypothesis of the MVT on the interval [-4,2]. sketch a graph of f. mx+a; when -4

OpenStudy (anonymous):

i know that the MVT says if f on [a,b] is continuous and differentiable, then there is an x=c within (a,b) such that f'(c)= f(b)-f(a) / b-a

OpenStudy (bahrom7893):

f'(c) = mx+a (first condition) a = -4, b = 2 (interval a to b) f(b)-f(a) / 2+4 = mx-4 f(b)-f(a)/6 = mx - 4

OpenStudy (bahrom7893):

or f(b)-f(a) = 6(mx - 4) hmm im kinda confused now

OpenStudy (anonymous):

i know the feeling

OpenStudy (anonymous):

do you know anyone on the site who's good at this kind of problem?

OpenStudy (bahrom7893):

@Zarkon @TuringTest can u guys take a look?

OpenStudy (turingtest):

I assume that is supposed to be ,x+a when \[-4\le x\le 2\]? (you have -2)

OpenStudy (turingtest):

mx+a *

OpenStudy (anonymous):

the negative 2 is correct

OpenStudy (turingtest):

Oh I misread sorry

OpenStudy (turingtest):

I get it now...

OpenStudy (zarkon):

this is more tedious than hard.

OpenStudy (zarkon):

you need to make f continuous and differentiable.

OpenStudy (zarkon):

looks like c and b need to be zero

OpenStudy (anonymous):

why? because they aren't (-)?

OpenStudy (zarkon):

to be continuous you need \[\lim_{x\to 0}f(x)=f(0)=0\]

OpenStudy (zarkon):

\[\lim_{x\to 0^{+}}f(x)=\lim_{x\to 0^{+}}(Mx+c)=M\cdot 0+c\] this must be 0 thus \(c=0\)

OpenStudy (anonymous):

ok. but then how do you find M? it could be anything in the interval, right?

OpenStudy (zarkon):

you get M from the differentiability condition.

OpenStudy (anonymous):

what does that mean?

OpenStudy (zarkon):

\[f'(0)=\lim_{h\to0}\frac{f(x+h)-f(x)}{h}\] must exist

OpenStudy (zarkon):

\[f'(0)=\lim_{h\to0}\frac{f(0+h)-f(0)}{h}\]

OpenStudy (turingtest):

does M=2 ? (I'm trying to follow along as well..)

OpenStudy (zarkon):

looking at the limit as h goes to zero from the right \[\lim_{h\to0^+}\frac{f(0+h)-f(0)}{h}=\lim_{h\to0^+}\frac{Mh-0}{h}=M\]

OpenStudy (zarkon):

\[\lim_{h\to0^-}\frac{f(0+h)-f(0)}{h}=\lim_{h\to0^+}\frac{h^2+2h-0}{h}=2\] thus, as TT stated, M=2

OpenStudy (anonymous):

ok, I need to start from a beginning. I'm lost as to how we got there. so if f(x)=x^2+2x+b when -2<x<0 first, i find f'(x) in the interval? f'(x)=2x+2, right? f'(x)=0 when x=-1?

OpenStudy (zarkon):

for get a - sign on the limit \[\lim_{h\to0^-}\frac{f(0+h)-f(0)}{h}=\lim_{h\to0^-}\frac{h^2+2h-0}{h}=2\]

OpenStudy (turingtest):

Thank you Zarkon for yet another brilliant explanation :D (sorry if I interrupted it, I just wanted to make sure I understood)

OpenStudy (anonymous):

not at all.

OpenStudy (zarkon):

it's all good TT :)

OpenStudy (anonymous):

I'm sorry, i looked at it again, and I'm more confused. I get how you find M, but not c

OpenStudy (turingtest):

what is the limit of f(x) as x goes to 0 from the right?

OpenStudy (anonymous):

0?

OpenStudy (turingtest):

and as x goes to zero from the left ?

OpenStudy (anonymous):

no holes, so it's 0, right?

OpenStudy (turingtest):

sorry no it's c

OpenStudy (anonymous):

so, is a=0 as well?

OpenStudy (zarkon):

remember what a typed above... \[\lim_{x\to 0^{+}}f(x)=\lim_{x\to 0^{+}}(Mx+c)=M\cdot 0+c=c\]

OpenStudy (anonymous):

Oook! so c is the limit.

OpenStudy (turingtest):

I messed up zarkons explanation, let me try again...

OpenStudy (turingtest):

what is the limit of f as x goes to zero from the left? (that is when f(x)=x^2+2x+b)

OpenStudy (anonymous):

is it b? or is it f'(x)? =2x+2?

OpenStudy (turingtest):

right, and the value of f(0) is defined to be...?

OpenStudy (anonymous):

b

OpenStudy (turingtest):

it's b...

OpenStudy (turingtest):

no f(0) is defined to be zero, yes?

OpenStudy (anonymous):

if x=0 f(0)=0^2 + (2*0) + b right? how do you get b=0?

OpenStudy (anonymous):

so, i need to find where f'(x)=0 for each condition? find all crotocal numbers for each interval?

OpenStudy (anonymous):

**critical**

OpenStudy (turingtest):

\[\lim_{x\to0^-}f(x)=b\]\[\lim_{x\to0^+}f(x)=c\]\[f(0)=0\]By definition thefunction will only be continuous at zero if\[\lim_{x\to0}f(x)=f(0)\]hence this is only continuous when\[\lim_{x\to0^-}f(x)=b=\lim_{x\to0^+}f(x)=c=f(0)=0\]

OpenStudy (anonymous):

ok. that makes sense!

OpenStudy (turingtest):

awesome, happy to help all credit to Zarkon though; this would've taken me ages to figure out alone!

OpenStudy (anonymous):

limx→0 f(x)=a

OpenStudy (anonymous):

is a the same?

OpenStudy (turingtest):

are we talking about the same problem? if so, then no...

OpenStudy (anonymous):

hmm. yea, i was talking about the same one. how do you find the lim of a?

OpenStudy (zarkon):

you need to look at f(-2) and \(\lim_{x\to -2^{-}}f(x)\)

OpenStudy (anonymous):

lim f(x) = lim(mx+a) = m⋅0+a=a x→0+ x→0+

OpenStudy (anonymous):

right? except with a negative. not a positive in the x-->0-

OpenStudy (turingtest):

no, f is not defined that way for x around 0

OpenStudy (zarkon):

we are not on the same page here...look at what i typed. above

OpenStudy (turingtest):

right, what is f defined as in the interval we are talking about?

OpenStudy (anonymous):

lim f(x) = lim f(mx+a) = m(-2) + a x→−2(-) x→−2(-)

OpenStudy (anonymous):

f(x)=mx+a in the interval, yes?

OpenStudy (turingtest):

from the left, yeah

OpenStudy (anonymous):

but the limit is -2, not 0. i got confused about the 0 because we just finished learning about critical #s and I keeo trying t set everything = 0

OpenStudy (anonymous):

**keep

OpenStudy (turingtest):

you always gotta look at the value that x is said to be approaching in the limit it could be anythin; x can go to infinity, zero, whatever

OpenStudy (anonymous):

so, we have f(x)=-2m + a and lim x->0 for f(x) is -2

OpenStudy (anonymous):

we find m using the definition of a derivative

OpenStudy (anonymous):

like you did with M

OpenStudy (anonymous):

but how do you get a?

OpenStudy (zarkon):

you will need to find m first

OpenStudy (turingtest):

derivative trick again

OpenStudy (turingtest):

I though he had found m

OpenStudy (zarkon):

he found M not m

OpenStudy (turingtest):

oh right...

OpenStudy (zarkon):

poor choice of variables :)

OpenStudy (turingtest):

so yeah, derivative trick again f'(-2) from the right must equal f'(-2) on the left

OpenStudy (anonymous):

how do you mean?

OpenStudy (turingtest):

same thing we did to find M (another reason it confused me)

OpenStudy (turingtest):

read above Zarkon's method if you didn't catch it

OpenStudy (anonymous):

yea, i'm just wndering if it'd be [f(m+h) - f(m)] / h

OpenStudy (turingtest):

it's always[f(x+h) - f(x)] / h

OpenStudy (anonymous):

do you sub the whole mx+a in for x here?

OpenStudy (turingtest):

f(x+h)=m(x+h)+a gotta remember these kinds of things ;)

OpenStudy (anonymous):

understanding them the first time around would've helped too, i guess...

OpenStudy (turingtest):

yeah, lol sometimes they don't teach first principles well :( anyway think you can get M ?

OpenStudy (anonymous):

so, you don't sub the function into the formula, you sub the formula into the function?

OpenStudy (anonymous):

(m(x+h)+a - mx+a) / h ?

OpenStudy (turingtest):

(m(x+h)+a - (mx+a) / h

OpenStudy (anonymous):

(mx+mh+a-mx-a) / h =mh / h

OpenStudy (anonymous):

would it be -2=mh / h?

OpenStudy (turingtest):

we are talking about M, yes?

OpenStudy (anonymous):

m=-2

OpenStudy (anonymous):

little m

OpenStudy (anonymous):

in the first condition

OpenStudy (turingtest):

oh right, so what about big M ?